OpenStudy (anonymous):

Rearrange this equation to isolate c a=b(1/c - 1/d)

OpenStudy (anonymous):

The formula is: \[a=b\left(\frac{1}{c}-\frac{1}{d}\right)\] Right?

OpenStudy (anonymous):

Yes

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

Okay. so: you have: \[\eqalign{ &1.\phantom{space}a=b\left(\frac{1}{c}-\frac{1}{d}\right) \\ &2.\phantom{space}\frac{a}{b}=\frac{1}{c}-\frac{1}{d} \\ &3.\phantom{space}\frac{a}{b}+\frac{1}{d}=\frac{1}{c} \\ &4.\phantom{space}c\left(\frac{a}{b}+\frac{1}{d}\right)=1 \\ &5.\phantom{space}c=\frac{1}{\frac{a}{b}+\frac{1}{d}} \\ }\] Cool? If there was something you didn't understand, I took the liberty of numbering each step :-D